Some time ago, one of the United States was facing a budget deficit of $1 billion. The population of the state was approximately 10.5 million. How much would each person in the state have to contribute to pay a $1 billion deficit? Round your answer to the nearest dollar.
Found 2 solutions by Boreal, MathTherapy:
Answer by Boreal(15235)

billion/number of people gives $/person, so that is the operation that needs to be done
10*10^8/1.05*10^7=9.523*10=$95.23
Note:
I can use 1*10^9/10.5*10^6 and get 0.09523*10^3, and moving the decimal 3 places to the right get the same answer. I prefer to have the constant in front of the 10 power larger in the numerator than in the denominator.

Answer by MathTherapy(10557)

Some time ago, one of the United States was facing a budget deficit of $1 billion. The population of the state was approximately 10.5 million. How much would each person in the state have to contribute to pay a $1 billion deficit? Round your answer to the nearest dollar.
All you have to do is divide 1,000 million (1 billion) by 10.5 million, or: 
That's ALL! Nothing else to it! Nothing COMPLEX &/or CONFUSING!